Find three consecutive odd numbers such that the difference of five times the third number and three times the first number is 54. Separate your answers with a comma. No space between all inputs.
Answer by checkley77(12844)

Let x, (x+2) & ( x+4) be the three numbers.
5(x+4)-3x=54
5x+20-3x=54
2x=54-20
2x=34
x=34/2
x=17 ans. for the smallest number.
17+4=21 ans. for the laegest number.
Proof:
5*21-3*17=54
105-51=54
54=54
(17,19,21) are the three numbers.